# Revealing Multimodal Causality with Large Language Models

### v1

@@ 4.4 Ablation Study
We conduct an ablation study to evaluate the effectiveness of each module in MLLM-CD. We compare the full MLLM-CD with the following variants: (1) MLLM-CD w/o Contrastive Factor Discovery (w/o CFD): This variant does not perform contrastive factor discovery, and instead, only uses a simple method, similar to [20], for factor identification. (2) MLLM-CD w/o Counterfactual Reasoning (w/o CR): This variant only uses the contrastive factor discovery module and the causal structure discovery module. (3) MLLM-CD w/o Both (w/o Both): This variant only uses a simple factor identification prompt from [20] and then performs causal structure discovery without any multimodal reasoning. The performance of these variants with Gemini 2.0 on two datasets is shown in Table 3. Please refer to Appendix D.6 for the full results. One can observe that excluding either the contrastive factor discovery or the counterfactual reasoning module leads to a significant drop in performance. In specific, contrastive factor discovery shows more impact on the accuracy and completeness of factor identification. The counterfactual reasoning module, on the other hand, shows more effectiveness in refining the causal structure, as we can see from the comparison of MLLM-CD w/o CR and the full MLLM-CD. Meanwhile, it contributes to larger improvements on the relatively smaller dataset (Lung Cancer) than the larger one (MAG). This indicates the importance of introducing plausible counterfactuals based on the MLLM’s knowledge and reasoning capabilities to alleviate the data scarcity issue and possible data noise. By combining both modules, they jointly contribute to the improvement of causal factor and structure discovery, demonstrating the effectiveness of MLLM-CD.
5 Conclusions
@@
This paper presents MLLM-CD, a novel framework for multimodal causal discovery that leverages the multimodal understanding and reasoning capabilities of MLLMs. MLLM-CD consists of three main components: (1) a contrastive factor discovery module to identify accurate and comprehensive causal factors from multimodal unstructured data in a novel multimodal contrastive manner, (2) a causal structure discovery module to learn the causal relationships among the identified factors, and (3) a multimodal counterfactual reasoning module to refine the discovered results by generating and validating counterfactual samples. We conduct extensive experiments on both synthetic and real-world multimodal datasets, demonstrating the superior performance in causal factor and structure discovery and the practical applicability of MLLM-CD. In future work, we will explore larger-scale multimodal unstructured CD. See Appendix B for more discussions on limitations and future work.
@@ Appendix A Table of Notations
$\mathcal{D}_{S}$ The structured multimodal dataset.
$\mathcal{D}_{S}^{(t)}$ The structured multimodal dataset at the $t$-th iteration.
$\mathcal{D}_{CF}^{(t)}$ The validated counterfactual samples at the $t$-th iteration.
@@
$\mathbf{V}_{\rm CFD}$ The set of factors identified by our CFD module.
$\mathbf{V}_{\rm direct}$ The set of factors identified by direct, non-contrastive methods.
@@
$\mathcal{G}^{*}_{\mathbf{V}}$ The true causal subgraph induced by $\mathbf{V}$.
@@
$\mathcal{G}_{\rm CFD}$ The causal graph inferred from $\mathbf{V}_{\rm CFD}$.
$\mathcal{G}_{\rm direct}$ The causal graph inferred from $\mathbf{V}_{\rm direct}$.
@@
$\mathbb{S}(\mathcal{G})$ The set of causal structures deemed compatible with the evidence from $\mathcal{D}_{S}$.
@@ Appendix C Algorithm
13:  Annotate values: $\mathcal{D}_{S}^{(t)}\leftarrow\Psi(\mathcal{D}^{(t)},\mathbf{V}^{(t)},\mathbf{p}_{\rm a})$.
@@
15:  Infer causal graph: $\mathcal{G}^{(t)}\leftarrow\mathcal{C}(\mathcal{D}_{S}^{(t)},\mathbf{V}^{(t)}\cup\{Y\})$.
@@
17:  Initialize set of validated counterfactuals $\mathcal{D}_{CF}^{(t)}\leftarrow\emptyset$.
@@
25:     $R_{\rm indep}(\mathbf{S}^{\prime}_{k},V_{a},\mathcal{G}^{(t)})\leftarrow\frac{\Sigma_{V_{j}\in{\rm NonDesc}(V_{a},\mathcal{G}^{(t)})}\mathbb{I}[\Delta v_{kj}\geq\epsilon]}{|V_{j}\in{\rm NonDesc}(V_{a},\mathcal{G}^{(t)})|}$.
@@
28:      $\mathcal{D}^{(t)}_{CF}\leftarrow\mathcal{D}^{(t)}_{CF}\cup\{\mathbf{X}^{\prime}_{k}\}$. $\triangleright$ Store multimodal counterfactual samples
@@
32:  $\mathcal{D}^{(t+1)}\leftarrow\mathcal{D}^{(t)}\cup\mathcal{D}^{(t)}_{CF}$. $\triangleright$ Augment multimodal dataset
33:end while
34:return $\mathbf{V}^{(T)},\mathcal{G}^{(T)}$.
@@ D.2 Details on Prompts
Prompt D.5: An Example of Factor Consolidation Prompt $\mathbf{p}_{\rm m}$ # Factor Deduplication and Refinement Below is a list of factors identified from different contrastive analyses: Factor 1: {factor 1} - 1: [Positive Criterion] - 0: [Netral Criterion] - -1: [Negative Criterion] Please: 1. Identify and merge similar factors; Each factor should be distinct and not overlap with others and be specific to one aspect; Avoid general factors like "overall quality" 2. Refine factor definitions for clarity and precision 3. Output a consolidated list of the most important and non-overlapping factors Only output the final results using this exact format for each factor: **Factor Name** - 1: [Positive Criterion] - 0: [Otherwise; or not mentioned] - -1: [Negative Criterion]
Prompt D.6: An Example of Annotation Prompt $\mathbf{p}_{\rm a}$ # Factor Annotation Please annotate the following samples based on these factors: ## Factor 1: {factor 1} - 1: [Positive Criterion] - 0: [Netral Criterion] - -1: [Negative Criterion] # Samples to Annotate ## Sample 1: - **Score**: {score 1} - **Review**: {review 1} - **Image**: {image 1} # Task: For each sample, assign a value (-1, 0, or 1) to each factor based on the criteria above. # Output Format Please format your response strictly as follows: **Sample [ID]**: - Factor 1 ([factor1 name]): [Value] - Factor 2 ([factor2 name]): [Value] Repeat for all samples.
@@ D.4 The Full Results on MAG Dataset
The full quantitative results for factor identification and structure discovery on the MAG dataset are shown in Table 5. Here, we also report the average results across all MLLM backbones. From these results, it is evident that MLLM-CD outperforms other baselines in most cases. META achieves a comparable performance in causal structure discovery, especially in terms of adjacency precision. This highlights the capability of MLLMs to identify potential causal relationships using their inherent knowledge, sometimes rivaling those derived from statistical methods. While COAT shows strong precision in causal factor discovery, it lacks comprehensiveness, which limits its overall effectiveness.
We also illustrate the visual comparison of causal graphs discovered by different methods and MLLM backbones on the MAG dataset. The results are shown in Figures 9, 10, 11 and 12. Several key observations can be made:
@@ D.5 The Full Results on Lung Cancer Dataset
Table 6: Causal factor identification and structure discovery performance on the Lung Cancer dataset.44footnotemark: 4
@@ Appendix E Broader Impacts
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  4. Justification: See the end of the Introduction section.
  5. Guidelines:
  6. 2.
    Limitations
  8. Answer: [Yes]
  9. Justification: See Section 5 and Appendix B.
  10. Guidelines:
  11. 3.
    Theory assumptions and proofs
  13. Answer: [Yes]
  14. Justification: See Appendix D.
  15. Guidelines:
  16. 4.
    Experimental result reproducibility
  18. Answer: [Yes]
  19. Justification: See Section 4.1 and Appendix E for detailed experimental settings and see Abstract for the link of implementation code and data.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
      1. (a)
      2. (b)
      3. (c)
      4. (d)
  21. 5.
    Open access to data and code
  23. Answer: [Yes]
  25. Guidelines:
  26. 6.
    Experimental setting/details
  28. Answer: [Yes]
  29. Justification: See Section 4.1 and Appendix E for detailed experimental settings.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
  31. 7.
    Experiment statistical significance
  33. Answer: [Yes]
  34. Justification: Standard deviations are included.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
  36. 8.
    Experiments compute resources
  38. Answer: [Yes]
  39. Justification: See Appendix E for details of experimental environment.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
  41. 9.
    Code of ethics
  43. Answer: [Yes]
  44. Justification: Authors follow the NeurIPS Code of Ethics.
  45. Guidelines:
  46. 10.
    Broader impacts
  48. Answer: [Yes]
  49. Justification: See Appendix F.
  50. Guidelines:
  51. 11.
    Safeguards
  53. Answer: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
  56. 12.
    Licenses for existing assets
  58. Answer: [Yes]
  59. Justification: See Appendix E for the license and terms of use of the datasets used in this paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
  61. 13.
    New assets
  63. Answer: [Yes]
  64. Justification: See Appendix E for the details of the constructed datasets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
  66. 14.
    Crowdsourcing and research with human subjects
  68. Answer: [N/A]
  70. Guidelines:
  71. 15.
  73. Answer: [N/A]
  75. Guidelines:
  76. 16.
    Declaration of LLM usage
  78. Answer: [N/A]
  80. Guidelines:

### v2

@@ 4.4 Ablation Study
We conduct an ablation study to evaluate the effectiveness of each module in MLLM-CD. We compare the full MLLM-CD with the following variants: (1) MLLM-CD w/o Contrastive Factor Discovery (w/o CFD): This variant does not perform contrastive factor discovery, and instead, only uses a simple method, similar to [20], for factor identification. (2) MLLM-CD w/o Counterfactual Reasoning (w/o CR): This variant only uses the contrastive factor discovery module and the causal structure discovery module. (3) MLLM-CD w/o Both (w/o Both): This variant only uses a simple factor identification prompt from [20] and then performs causal structure discovery without any multimodal reasoning. The performance of these variants with Gemini 2.0 on two datasets is shown in Table 3.
@@
Please refer to Appendix D.6 for the full results. One can observe that excluding either the contrastive factor discovery or the counterfactual reasoning module leads to a significant drop in performance. In specific, contrastive factor discovery shows more impact on the accuracy and completeness of factor identification. The counterfactual reasoning module, on the other hand, shows more effectiveness in refining the causal structure, as we can see from the comparison of MLLM-CD w/o CR and the full MLLM-CD. Meanwhile, it contributes to larger improvements on the relatively smaller dataset (Lung Cancer) than the larger one (MAG). This indicates the importance of introducing plausible counterfactuals based on the MLLM’s knowledge and reasoning capabilities to alleviate the data scarcity issue and possible data noise. By combining both modules, they jointly contribute to the improvement of causal factor and structure discovery, demonstrating the effectiveness of MLLM-CD.
Table 4: Ablation Study of the CFD module on the MAG and Lung Cancer datasets.
Dataset Variant NP $\uparrow$ NR $\uparrow$ NF $\uparrow$
MAG Random $\bf 0.94\pm 0.10$ $0.60\pm 0.06$ $0.73\pm 0.07$
Simple Pair $0.79\pm 0.14$ $0.70\pm 0.13$ $0.75\pm 0.14$
Intra-modal $0.88\pm 0.13$ $0.78\pm 0.11$ $0.82\pm 0.12$
Inter-modal $0.67\pm 0.19$ $0.59\pm 0.17$ $0.63\pm 0.18$
MLLM-CD $0.86\pm 0.05$ $\bf 0.89\pm 0.00$ $\bf 0.87\pm 0.03$
Lung Random $0.65\pm 0.09$ $0.60\pm 0.00$ $0.62\pm 0.04$
Simple Pair $0.50\pm 0.00$ $0.60\pm 0.00$ $0.55\pm 0.00$
Intra-modal $0.80\pm 0.00$ $0.80\pm 0.00$ $0.80\pm 0.00$
Inter-modal $0.53\pm 0.12$ $0.53\pm 0.12$ $0.53\pm 0.12$
MLLM-CD $\bf 0.94\pm 0.10$ $\bf 1.00\pm 0.00$ $\bf 0.97\pm 0.05$
In addition, we further evaluate the effectiveness of different sampling strategies used in our contrastive factor discovery module. We compare the following variants: (1) Random Sampling (Random): This variant randomly samples data points from each category of the target variable. (2) Simple Pairwise Sampling (Simple Pair): This variant pairs samples with the largest difference in the target variable. (3) Intra-modal Contrastive Sampling (Intra-modal): This variant only considers intra-modal contrastive pairs within each modality. (4) Inter-modal Contrastive Sampling (Inter-modal): This variant only considers inter-modal contrastive pairs across different modalities for sampling. The key quantitative results are shown in Table 4, and we have the following observations: (1) Incorporating finer-grained signals by finding semantic-level contrastive pairs within each modality, the intra-modal contrastive sampling further improves the completeness of factor discovery. However, it often overlooks the cross-modal interactions. (2) Inter-modal contrastive sampling acts as an ideal complement to intra-modal strategy, as it examines across modalities and can effectively identify factors missed by intra-modal. Using inter-modal only has lower quantitative results, because it is prompted to emphasize factors tied to cross-modal interactions, as a complement to intra-modal strategy, rather than aiming to discover all factors. (3) By combining both intra-modal and inter-modal strategies, MLLM-CD achieves the best overall performance in factor discovery, demonstrating the effectiveness of our contrastive factor discovery module.
5 Conclusion
@@
This paper presents MLLM-CD, a novel framework for multimodal causal discovery that leverages the multimodal understanding and reasoning capabilities of MLLMs. MLLM-CD consists of three main components: (1) a contrastive factor discovery module to identify accurate and comprehensive causal factors from multimodal unstructured data in a novel multimodal contrastive manner, (2) a causal structure discovery module to learn the causal relationships among the identified factors, and (3) a multimodal counterfactual reasoning module to refine the discovered results by generating and validating counterfactual samples. We conduct extensive experiments on both synthetic and real-world multimodal datasets, demonstrating the superior performance in causal factor and structure discovery and the practical applicability of MLLM-CD.
Limitations. While promising, our approach faces several limitations. A primary challenge is the scarcity of large-scale, diverse benchmark datasets for multimodal causal discovery. While the datasets used in this work are thoughtfully curated, they are constrained by small sample sizes, and their ground truth causal graphs depend on domain expert knowledge. Secondly, the range of modalities MLLM-CD can effectively process is bound by the capabilities of the core MLLM, presenting challenges for integrating specialized modalities like sensor data or genomic sequences. Finally, the overall effectiveness of the proposed framework is dependent on the sophisticated reasoning and generation capabilities of MLLMs, which may still exhibit hallucinations, reflect biases from their training data, or fail to capture subtle causal nuances, despite recent advancements in model alignment.
Future Work. These limitations open several avenues for future research. A crucial direction is the development of larger, more diverse, and rigorously annotated benchmark datasets, potentially using semi-automated annotation or simulation-based approaches, which will also allow for exploring the scalability of MLLM-CD. We will also focus on enhancing the framework to handle a wider range of modalities by developing more sophisticated multimodal fusion and cross-modal reasoning mechanisms. To address the reliance on MLLM capabilities, we plan to develop methods to rigorously validate and improve MLLM outputs within the causal discovery pipeline. This includes integrating with interventional analysis, designing uncertainty quantification for generations, and exploring techniques to mitigate biases and hallucinations specifically for causal reasoning.
Please refer to Appendix B for more discussions on limitations and future work.
Acknowledgments and Disclosure of Funding
We sincerely appreciate the reviewers’ valuable feedback. This work is partially supported by Australia ARC LP220100453 and ARC DP240100955. TLL is partially supported by the following Australian Research Council projects: FT220100318, DP220102121, LP220100527, LP220200949.
NeurIPS Paper Checklist
  1. 1.
    Claims
  3. Answer: [Yes]
  4. Justification: See the end of the Introduction section.
  5. Guidelines:
  6. 2.
    Limitations
  8. Answer: [Yes]
  9. Justification: See Section 5 and Appendix B.
  10. Guidelines:
  11. 3.
    Theory assumptions and proofs
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical results.
  15. Guidelines:
  16. 4.
    Experimental result reproducibility
  18. Answer: [Yes]
  19. Justification: See Section 4.1 and Appendix D for detailed experimental settings and see Abstract for the link of implementation code and data.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
      1. (a)
      2. (b)
      3. (c)
      4. (d)
  21. 5.
    Open access to data and code
  23. Answer: [Yes]
  25. Guidelines:
  26. 6.
    Experimental setting/details
  28. Answer: [Yes]
  29. Justification: See Section 4.1 and Appendix D for detailed experimental settings.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
  31. 7.
    Experiment statistical significance
  33. Answer: [Yes]
  34. Justification: Standard deviations are included.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
  36. 8.
    Experiments compute resources
  38. Answer: [Yes]
  39. Justification: See Appendix D for details of experimental environment.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
  41. 9.
    Code of ethics
  43. Answer: [Yes]
  44. Justification: Authors follow the NeurIPS Code of Ethics.
  45. Guidelines:
  46. 10.
    Broader impacts
  48. Answer: [Yes]
  49. Justification: See Appendix E.
  50. Guidelines:
  51. 11.
    Safeguards
  53. Answer: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
  56. 12.
    Licenses for existing assets
  58. Answer: [Yes]
  59. Justification: See Appendix D for the license and terms of use of the datasets used in this paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
  61. 13.
    New assets
  63. Answer: [Yes]
  64. Justification: See Appendix D for the details of the constructed datasets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
  66. 14.
    Crowdsourcing and research with human subjects
  68. Answer: [N/A]
  70. Guidelines:
  71. 15.
  73. Answer: [N/A]
  75. Guidelines:
  76. 16.
    Declaration of LLM usage
  78. Answer: [N/A]
  80. Guidelines:
@@ Appendix A Table of Notations
$\mathcal{D}_{\rm S}$ The structured multimodal dataset.
$\mathcal{D}_{\rm S}^{(t)}$ The structured multimodal dataset at the $t$-th iteration.
$\mathcal{D}_{\rm CF}^{(t)}$ The validated counterfactual samples at the $t$-th iteration.
@@ Appendix C Algorithm
13:  Annotate values: $\mathcal{D}_{\rm S}^{(t)}\leftarrow\Psi(\mathcal{D}^{(t)},\mathbf{V}^{(t)},\mathbf{p}_{\rm a})$.
@@
15:  Infer causal graph: $\mathcal{G}^{(t)}\leftarrow\mathcal{C}(\mathcal{D}_{\rm S}^{(t)},\mathbf{V}^{(t)}\cup\{Y\})$.
@@
17:  Initialize set of validated counterfactuals $\mathcal{D}_{\rm CF}^{(t)}\leftarrow\emptyset$.
@@
25:     $R_{\rm indep}(\mathbf{S}^{\prime}_{k},V_{a},\mathcal{G}^{(t)})\leftarrow\frac{\Sigma_{V_{j}\in{\rm NonDesc}(V_{a},\mathcal{G}^{(t)})}\mathbb{I}[\Delta v_{kj}\geq\epsilon]}{|{\rm NonDesc}(V_{a},\mathcal{G}^{(t)})|}$.
@@
28:      $\mathcal{D}^{(t)}_{\rm CF}\leftarrow\mathcal{D}^{(t)}_{\rm CF}\cup\{\mathbf{X}^{\prime}_{k}\}$. $\triangleright$ Store multimodal counterfactual samples
@@
32:  $\mathcal{D}^{(t+1)}\leftarrow\mathcal{D}^{(t)}\cup\mathcal{D}^{(t)}_{\rm CF}$. $\triangleright$ Augment multimodal dataset
33:  $t\leftarrow t+1$.
34:end while
35:return $\mathbf{V}^{(T)},\mathcal{G}^{(T)}$.
@@ D.2 Details on Prompts
Prompt D.5: An Example of Factor Consolidation Prompt $\mathbf{p}_{\rm m}$ # Factor Deduplication and Refinement Below is a list of factors identified from different contrastive analyses: Factor 1: {factor 1} - 1: [Positive Criterion] - 0: [Neutral Criterion] - -1: [Negative Criterion] Please: 1. Identify and merge similar factors; Each factor should be distinct and not overlap with others and be specific to one aspect; Avoid general factors like "overall quality" 2. Refine factor definitions for clarity and precision 3. Output a consolidated list of the most important and non-overlapping factors Only output the final results using this exact format for each factor: **Factor Name** - 1: [Positive Criterion] - 0: [Otherwise; or not mentioned] - -1: [Negative Criterion]
Prompt D.6: An Example of Annotation Prompt $\mathbf{p}_{\rm a}$ # Factor Annotation Please annotate the following samples based on these factors: ## Factor 1: {factor 1} - 1: [Positive Criterion] - 0: [Neutral Criterion] - -1: [Negative Criterion] # Samples to Annotate ## Sample 1: - **Score**: {score 1} - **Review**: {review 1} - **Image**: {image 1} # Task: For each sample, assign a value (-1, 0, or 1) to each factor based on the criteria above. # Output Format Please format your response strictly as follows: **Sample [ID]**: - Factor 1 ([factor1 name]): [Value] - Factor 2 ([factor2 name]): [Value] Repeat for all samples.
@@ D.4 The Full Results on MAG Dataset
The full quantitative results for factor identification and structure discovery on the MAG dataset are shown in Table 6. Here, we also illustrate the visual comparison of causal graphs discovered by different methods and MLLM backbones on the MAG dataset. The results are shown in Figures 9, 10, 11 and 12. Several key observations can be made:
@@ D.5 The Full Results on Lung Cancer Dataset
Table 7: Causal factor identification and structure discovery performance on the Lung Cancer dataset.77footnotemark: 7